Developing Hypotheses about how Capacitors Combine
Answer #1

Capacitors in parallel combination


   capacitors are used to store some electrical energy, and the capacitance of the capacitors connected in parallel is

we know the relation Q = C*V

   now if two capacitors connected in parallel of capacitance C1,C2


let the total charge is Qt = Q1+Q2

and the voltage across each cpacitor is same so

   Ct.V = C1*V +C2*V

   Ct = C1 + C2

and the capacitance of a capacitor is C = epsilon not *A/d epsilon not is permitivity of free space


               Ct = C1+C2
               Ct = epsilon not/d (A1+A2)

               Ct = epsilon not A_eff/d


so the effective area A is A_eff = A1+A2